OpenStudy (anonymous):

A container states that it should be stored at a temperature below 35c . Convert the temperature to degrees Fahrenheit. Round to the nearest whole number, if necessary.

OpenStudy (phi):

do you know the formula?

OpenStudy (phi):

c= 5/9(f-32)

OpenStudy (phi):

I think you should replace the c with 35, and solve for f. 1st step, mult both sides by 9/5

OpenStudy (anonymous):

35/1 * 9/5 = 315/5 = 63/1

OpenStudy (phi):

so you have 63= f-32 now what? you can check your answer by typing 35 c in f= in the google search window

OpenStudy (anonymous):

63 isn't in the multiple choice:

OpenStudy (phi):

that is because you have not finished. You started with 35= (5/9)(f-32) 63= f-32 solve for f

OpenStudy (anonymous):

95

OpenStudy (phi):

if you don't know, add 32 to both sides.

OpenStudy (anonymous):

95

OpenStudy (anonymous):

thx

OpenStudy (phi):

you can check your answer by typing 35 c in f= in the google search window but yes. 35 c= 95 f

OpenStudy (anonymous):

i got another question what is a rotational symmetry?

OpenStudy (phi):

it means an object looks the same after you have rotated it by something other than 360º Example: a square. you can rotate it by 90º and it looks the same

OpenStudy (anonymous):

how about a triangle no right?

OpenStudy (phi):

an equilateral triangle, but not any old triangle

OpenStudy (anonymous):

so an equilateral triangle doesnt have a rotational symmetry

OpenStudy (anonymous):

how about a rhombus, crescent, do they have a rotational symmetry?

OpenStudy (phi):

equilateral triangle does have rotational symmetry rotate it 120º and it looks the same. 

OpenStudy (anonymous):

how about a crescent, rhombus

OpenStudy (phi):